OpenStudy (piercetheveil47):

Compare 6 times 10^8 to 3 times 10^6. 6 times 10^8 is 2,000 times larger than 3 times 10^6. 6 times 10^8 is 200 times larger than 3 times 10^6. 6 Times 10^8 is 20 times larger than 3 times 10^6. 6 Times 10^8 is 2 times larger than 3 Times 10^6.

OpenStudy (anonymous):

can I ask you something?

OpenStudy (piercetheveil47):

yes

OpenStudy (anonymous):

is that you in the picture?

OpenStudy (piercetheveil47):

no lmao :)

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

oh lol because I found tht=at picture on google

OpenStudy (piercetheveil47):

yeah xD

OpenStudy (anonymous):

oh..lol ?

OpenStudy (anonymous):

why did you put that as your profile pic? .-.

OpenStudy (piercetheveil47):

bc she's my fav model

OpenStudy (piercetheveil47):

but someone pleaaaseee help me w this question

OpenStudy (anonymous):

oh :)

OpenStudy (piercetheveil47):

@perl

OpenStudy (lyrae):

Divide the first number by the second.

OpenStudy (lyrae):

\[\frac{ 6 \times 10^8 }{ 3 \times10^6 }\]

OpenStudy (piercetheveil47):

is it A?

OpenStudy (lyrae):

Nope. What's 10^8/10^6 ?

OpenStudy (lyrae):

\[\frac{ x^a }{ x^b } = x^{a-b}\]

OpenStudy (piercetheveil47):

its confusing since its the scientific notation

OpenStudy (lyrae):

Let's try and substitute the numbers in to the formula I gave you. \[\frac{ 10^8 }{ 10^6 } = 10^{8-6} = 10^2\]What's 10^2 ?

OpenStudy (anonymous):

1,000

OpenStudy (lyrae):

@joker123 nah. 10^2 = 10*10. Does 10*10 = 1000?

OpenStudy (piercetheveil47):

100

OpenStudy (piercetheveil47):

but is that the scientific notation?

OpenStudy (anonymous):

Ummm. im in the 8th grade and it looks like scientific notation too me

OpenStudy (anonymous):

no 10 with the power of two equal 1000 10x10=100x10=1000

OpenStudy (piercetheveil47):

yeah i think so too

OpenStudy (piercetheveil47):

bc its the scientific notation not exponents

OpenStudy (piercetheveil47):

brush does anyone know the answer :(

OpenStudy (piercetheveil47):

*bruh

OpenStudy (anonymous):

I'm working on it

OpenStudy (lyrae):

Remeber what we began with \[\frac{ 6 \times 10^8 }{ 3 \times 10^6 }\] You divide this calculation into two parts:\[\frac{ 6 }{ 3 } \times \frac{ 10^8 }{ 10^6 }\] We did the second part, so we have \[\frac{ 6 }{ 3 } \times 10^2\] How do you continue from there?

OpenStudy (anonymous):

My mind is like WHAT!!!!!!!!!??????????????????????????

OpenStudy (piercetheveil47):

same

OpenStudy (piercetheveil47):

but i gtg

OpenStudy (anonymous):

Okay sorry I cant help I got stuck